OpenStudy (anonymous):

Simplify the following expression: (x − 8y) - (2x − 4y). If the final answer is written in the form Ax + By, what is the value of B?

OpenStudy (e.cociuba):

x-8y-2x+4y = -x-4y A=-1,B=-4

OpenStudy (anonymous):

thank you

OpenStudy (e.cociuba):

No problemo! :)

OpenStudy (anonymous):

thanks for the medal (:

OpenStudy (e.cociuba):

Np :)